OpenStudy (anonymous):

5(6-4r)≤-2(r-7) A) r≥8/7 B) r≤-8/7 C) r≥-8/7 D) r≤8/7

OpenStudy (danjs):

\[30 - 20r \le -2r + 14\]

OpenStudy (danjs):

\[30 -14 \le (-2+20)r\]

OpenStudy (danjs):

\[16 \le 18r\] \[r \ge \frac{ 16 }{ 18 }\] \[r \ge \frac{ 8 }{ 9 }\]

OpenStudy (anonymous):

thanks man

OpenStudy (danjs):

not sure if you mistyped the equation or the answers, but that's the solution to the inequality you gave

OpenStudy (anonymous):

oh yeah i spelled wrong 5(6-4r)≤-2(3r-7)*

OpenStudy (anonymous):

5(6-4r)≤-2(3r-7)*

OpenStudy (danjs):

oh

OpenStudy (anonymous):

sorry

OpenStudy (danjs):

\[30 - 20r \le -6r +14\]

OpenStudy (danjs):

\[30 - 14 \le (-6+20)r\]

OpenStudy (danjs):

\[16 \le 14 r\] \[r \ge \frac{ 16 }{ 14 }\] \[r \ge \frac{ 8 }{ 7 }\]

OpenStudy (anonymous):

thanks man